Ayana has $2.35 in nickels and dimes. If she has 33 coins in all, find the number of nickels and dimes


1
Expert's answer
2022-02-17T13:03:28-0500

let

x = the number of dimes, x is positive integer

y = the number of nickles, y is positive integer

1 dime = 10 cents

1 nickle = 5 cents

she has 33 coins in all

x + y = 33...........(1)


Ayana has $2.35 in nickels and dimes that is 235 cents


10x + 5y = 235 divide both sides by 5


2x + y = 47.............(2)


Subtracting equation 1 from equation 2, we get,

x=47-33=14

From equation (2)

y=47-2(14)=47-28=19

Therefore, there are 14 dimes and 19 nickels.